OpenStudy (anonymous):

Find the inverse of each: f(x)=4x-1

OpenStudy (anonymous):

\[y=x^2+5\]

OpenStudy (anonymous):

let y = 4x + 1 exchange x and y x = 4y + 1 4y = x - 1 y = (x - 1)/4 = (1/4)x - (1/4) f^-1(x) = (1/4)x - (1/4)